OpenStudy (anonymous):

the surface of a baseball is 177 cm^2. What is the diameter of the baseball

OpenStudy (campbell_st):

well the surface area of a sphere can be found using \[SA = 4 \pi r^2\] so substitute 177 and solve for r \[177 = 4 \pi r^2\]

OpenStudy (anonymous):

I'm not getting the answer

OpenStudy (campbell_st):

so what are you getting...?

OpenStudy (anonymous):

14.01

OpenStudy (anonymous):

To find 'r' of \[SA = 4 \pi r^2\] You rearrange the equation \[r = \sqrt{\frac{ SA }{ 4 \pi }}\] Then do what campbell_st said by substituting SA with 177

OpenStudy (anonymous):

The amount you figure out here would be the radius which you would then double to get the diameter